### v1

# Krylov Complexity for Open Quantum System: Dissipation and Decoherence

Abstract
We investigate Krylov complexity in open quantum systems using Lindblad master equations for bosonic bath models, with particular emphasis on the Caldeira–Leggett model. Krylov complexity is computed from the moments of the two-point function within the standard master equation framework. For the damped harmonic oscillator, the results reveal clear dissipative features in Krylov complexity. In the Caldeira–Leggett model, in the high-temperature limit, we find that Krylov complexity saturates in the full system and reproduces the expected dissipative behavior when the decoherence term is suppressed in the master equation. Conversely, when the dissipative term is suppressed, the contribution from decoherence exhibits the familiar oscillatory dynamics of the coherent system, along with additional novel features. However, Krylov complexity appears insensitive to the onset of decoherence, as no clear distinctive signature is observed. We attribute this to the fact that Krylov complexity is defined in the Krylov basis, which does not coincide with the conventional basis typically used to study decoherence.
1 Introduction
Understanding the mechanisms of dissipation Ferialdi_2017 ; Ramazanoglu_2009 and decoherence Schlosshauer_2019 ; Habib:1998ai in open quantum systems is a fundamental challenge in physics, with far-reaching implications for fields ranging from low-temperature phenomena and early-universe cosmology to modern applications like quantum computing horodecki2013fundamental ; Skrzypczyk_2014 ; Lesgourgues:1996jc ; PhysRevA.52.R2493 ; Aharonov_2000 ; Ashida:2020dkc . Decoherence occurs when a quantum system interacts with its environment, leading to a loss of phase coherence. For a comprehensive review of this subject, interested readers are referred to Schlosshauer:2019ewh ; Zurek:2003zz . Dissipation, in contrast, involves the irreversible transfer of energy from the system to its environment, often resulting in the damping of quantum dynamics. A useful quantity for characterizing dissipation in a quantum system is the probability amplitude Xu:2018bhd , while the degree of decoherence—or, equivalently, the deviation of an open quantum system from a pure state—can be quantified using the decoherence function Zurek:2003zz ; open .
In recent years, quantum circuit complexity Susskind:2014moa ; Jefferson:2017sdb ; Nielsen:2005mkt has gained momentum as a powerful tool for exploring the properties and behavior of quantum systems, capturing the interest of researchers across multiple disciplines. The concept first rose to prominence in the context of the AdS/CFT correspondence and black hole physics, where it was developed as an extension of entanglement entropy Ryu:2006bv to probe the interior of a black hole Susskind:2014moa ; Susskind:2014rva ; Brown:2015bva ; Brown:2015lvg ; Carmi:2016wjl . However, its foundations were already well-established in quantum information theory, where it has served as a measure of the computational resources required to prepare or simulate quantum states Baiguera:2025dkc . Today, circuit complexity is widely applied to diverse quantum systems Jefferson:2017sdb ; Hackl:2018ptj ; Khan:2018rzm ; Chapman:2018hou ; Ali:2018fcz ; Ali:2018aon ; Ali:2019zcj , including condensed matter physics Liu_2020 ; Camargo_2019 ; Craps:2023rur , cosmology Bhattacharyya:2020rpy ; Bhattacharyya:2020kgu ; Haque:2021kdm , astrophysics Dixit:2023fke ; Jha:2025tde and quantum computing Baiguera:2025dkc . Interested readers are referred to these reviews Chapman:2021jbh ; Bhattacharyya:2021cwf ; Baiguera:2025dkc for more details.
Recent studies have shown that circuit complexity is sensitive to decoherence Bhattacharyya:2021fii ; Bhattacharyya_2023 ; Bhattacharyya_2024 ; Bhattacharyya:2025cxv ; Haque:2021hyw , a process that disrupts quantum coherence and affects the performance of quantum systems in various contexts Hornberger ; Schlosshauer_2019 ; BRE02 . This sensitivity makes circuit complexity a valuable diagnostic tool for understanding how environmental interactions impact quantum systems. More recently, an alternative measure, known as Krylov complexity, has been proposed as a complementary approach to quantifying the complexity of quantum dynamics Parker:2018yvk ; Barbon:2019wsy ; Avdoshkin:2019trj ; PhysRevB.102.085137 ; Balasubramanian:2022tpr . Unlike circuit complexity, which counts the minimum number of quantum gates needed to construct a state, Krylov complexity focuses on the evolution of operators in the Krylov space Parker:2018yvk , offering a unique perspective on the dynamical behavior of quantum systems.
Given the sensitivity of circuit complexity to decoherence, it is natural to explore whether Krylov complexity exhibits similar behavior, particularly in the presence of both decoherence and dissipation—two fundamental processes that degrade quantum information in open quantum systems. Decoherence arises from interactions with an external environment, causing a loss of quantum coherence, while dissipation involves energy exchange, often leading to the damping of quantum states. Determining whether Krylov complexity can serve as a robust probe of these phenomena is of great interest, as it could offer new insights into the behavior of complex quantum systems under realistic, non-ideal conditions.
We investigate the sensitivity of Krylov complexity to decoherence and dissipation, with the aim of assessing its usefulness as a reliable measure for studying these processes. Focusing on quantum systems such as the damped harmonic oscillator and the Caldeira–Leggett model, we explore the potential of Krylov complexity as a diagnostic tool. We further compare its performance with circuit complexity and evaluate its ability to capture the effects of environmental interactions. Our results have implications for the study of open quantum systems and contribute to the broader understanding of complexity in quantum mechanics.
Krylov complexity is defined in the Krylov subspace, originally developed to address large-scale linear algebra problems doi:10.1137/1036123 . It offers a novel approach to studying operator growth, with a key advantage being its ability to reduce high-dimensional systems to a one dimensional chain problem. Krylov complexity has proven particularly useful in diagnosing quantum phase transitions. In maximally chaotic systems, complexity grows at most exponentially Parker_2019 ; Barbon:2019wsy ; Rabinovici_2022 ; Rabinovici_2021 ; Bhattacharyya:2023dhp , whereas in open quantum systems it typically exhibits rapid initial growth followed by late-time saturation, with a maximum value lower than that of the corresponding closed system Bhattacharya_2023 ; Liu_2023 ; Bhattacharyya:2023grv . For more applications of Krylov complexity, readers are referred to these reviews Nandy_2025 ; Rabinovici:2025otw and to some recent works where it is studied for non-Hermitian systems Bhattacharya_2022 ; Bhattacharya_2023 ; Liu_2023 ; Bhattacharyya:2023grv ; Bhattacharjee:2023uwx ; NSSrivatsa:2023pby ; Bhattacharya:2023yec ; Bhattacharya:2024hto ; Chakrabarti:2025hsb ; Medina-Guerra:2025wxg ; Medina-Guerra:2025rwa ; Baggioli:2025knt , which motivated the study presented in this paper.
In this study, we apply Krylov complexity to investigate features of open quantum systems, focusing on systems exhibiting dissipative and decoherent effects, such as the damped harmonic oscillator and the Caldeira-Leggett model Caldeira:1982iu ; Bhattacharyya_2023 . The Caldeira-Leggett model describes a system coupled to a thermal bath of non-interacting harmonic oscillators, with dynamics governed by a Markovian master equation. This equation captures the slow, systematic evolution of the system compared to the environment, a hallmark of quantum Brownian motion.
The paper is organized as follows. In Sec. (2), we outline the details of Krylov complexity and its computation from the moments of the two-point function. In Sec. (3), we discuss the damped harmonic oscillator model and present its Krylov complexity results. In Sec. (4), we introduce the Caldeira-Leggett master equation, its assumptions, and its features, and analyze the associated Krylov complexity and its implications. Finally, in Sec. (5), we summarize our findings and discuss their broader implications for future research. In the Appendix (A), we have provided some additional details.
2 Krylov Complexity
In this paper, we investigate the growth of Krylov complexity in open quantum systems. We begin with a brief review of Krylov complexity, followed by an overview of the moments method used to compute it Parker_2019 ; dymarsky2020quantum ; Dymarsky_2021 ; vish123 . Krylov complexity is particularly valuable for its ability to project high-dimensional problems onto a lower-dimensional Krylov subspace. Our analysis starts in the Heisenberg picture of operator evolution, where time evolution naturally introduces the Liouvillian superoperator and leads to the construction of the Krylov basis. To understand this framework, consider the Heisenberg equation in the operator picture:
$\partial_{t}O(t)=i[H,O(t)],\quad O(t)=e^{-iHt}O(0)e^{iHt}.$ (1)
Using the Baker-Campbell-Hausdorff (BCH) formula, we can expand the operator evolution as:
$O(t)$ $=O+(it)[H,O]+\frac{(it)^{2}}{2!}[H,[H,O]]+\frac{(it)^{3}}{3!}[H,[H,[H,O]]]+\dots\,$,
$=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\mathcal{L}^{n}O$
$=e^{i\mathcal{L}t}O$ (2)
where $\mathcal{L}^{n}$ denotes nested applications of the Liouvillian superoperator, defined as the commutator $H,\cdot]$. The Krylov basis is defined through the nested application of $\mathcal{L}$ to the operator $O$. The Krylov subspace, within which the operator evolves, is spanned by the set of nested commutators:
$\mathrm{span}\{O,\mathcal{L}O,\mathcal{L}^{2}O,\mathcal{L}^{3}O,\dots,\mathcal{L}^{n}O\}.$ (3)
In the Krylov basis, the operator evolution is expressed as:
$O(t)=\sum_{n=0}^{\infty}i^{n}\phi_{n}(t)O_{n},$ (4)
where $O_{n}$ are the basis operators. The Krylov operator complexity is defined as Parker_2019 :
$K_{O}(t)=\sum_{n=0}^{\infty}n|\phi_{n}(t)|^{2},$ (5)
where $\phi_{n}(t)$ are the Krylov wavefunctions, representing the amplitude of the operator at site $n$ along the one-dimensional Krylov chain. Thus, Krylov complexity is the expectation value of the operator’s position along this chain. These wavefunctions are obtained by solving the Schrödinger-like equation Parker_2019 ; dymarsky2020quantum ; Dymarsky_2021 :
$\partial_{t}\phi_{n}(t)=b_{n}\phi_{n-1}(t)-b_{n+1}\phi_{n+1}(t)-a_{n}\phi_{n}(t),$ (6)
where $a_{n}$ and $b_{n}$ are the Lanczos coefficients. From (4) and (6), one can see that the system’s dynamics are fully encoded in the Lanczos coefficients. For unitary (closed) systems, the $a_{n}$ coefficients vanish. The Lanczos coefficients can be computed using methods such as the Lanczos algorithm or the moments of the autocorrelation function Parker_2019 ; dymarsky2020quantum ; Dymarsky_2021 ; vish123 .
Moments Method
The moments method involves computing the Lanczos coefficients via the two-point correlation function, also known as the autocorrelation function or return amplitude Dymarsky_2021 ; Balasubramanian:2022tpr :
$C(t)=\langle O(t)|O(0)\rangle=\phi_{0}(t).$ (7)
The moments are obtained by taking the $n$-th time derivative of the correlation function :
$\mu_{n}=\frac{d^{n}}{dt^{n}}\langle O(t)|O(0)\rangle.$ (8)
The Lanczos coefficients $\{b_{n}\}$ and $\{a_{n}\}$ are related to the moments $\{\mu_{n}\}$ via Hashimoto_2023 ; Parker_2019 ; dymarsky2020quantum ; Dymarsky_2021 ; vish123 :
$b^{2}_{n}(t)$ $=\frac{\tau_{n+1}(t)\tau_{n-1}(t)}{\tau_{n}^{2}(t)}\quad n\geq 0,$ (9)
$a_{n}(t)$ $=\frac{d}{dt}\log\left(\frac{\tau_{n}(t)}{\tau_{n-1}(t)}\right)\quad n\geq 0,$ (10)
where $\tau_{n}(t)$ are determinants of matrices constructed from the moments:
$\tau_{n}(t)=\begin{vmatrix}\mu_{0}&\mu_{1}&\mu_{2}&\cdots&\mu_{n}\\ \mu_{1}&\mu_{2}&\mu_{3}&\cdots&\mu_{n+1}\\ \mu_{2}&\mu_{3}&\mu_{4}&\cdots&\mu_{n+2}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n}&\mu_{n+1}&\mu_{n+2}&\cdots&\mu_{2n}\end{vmatrix}$, (11)
with $\tau_{-1}(t)=0$. Typically, the coefficients are evaluated at a cutoff time $t_{c}=0$, except in certain cases, such as conformal field theories (CFTs). The evolution of the operator, driven by the Liouvillian as shown in (2), is central to studying Krylov complexity, which quantifies operator growth. This can be visualized as follows:
[…]
Krylov Complexity for Open Quantum System
For open quantum systems governed by a Lindbladian $\mathcal{L}$, which describes non-unitary evolution due to environmental coupling, the standard Lanczos algorithm for Hermitian operators is insufficient. Instead, the bi-Lanczos algorithm is required to handle the non-Hermitian nature of the Lindbladian Bhattacharya_2023 ; Chakrabarti:2025hsb . This algorithm constructs a bi-orthogonal Krylov basis, yielding a tridiagonal matrix representation with complex diagonal elements $a_{n}$ and off-diagonal elements $b_{n}$ (superdiagonal) and $c_{n}$ (subdiagonal). These off-diagonals are generally complex and unequal ($b_{n}\neq c_{n}^{*}$), reflecting the dissipative dynamics. In moment-based methods for computing Krylov complexity, the product $b_{n}c_{n}$ often appears, but its complex nature can complicate calculations, especially when taking the square root for complexity metrics. A similarity transformation can symmetrize the off-diagonals to $\sqrt{b_{n}c_{n}}$, simplifying moment computations while preserving the spectrum.
The Lindbladian’s tridiagonal form in the bi-Lanczos basis is Bhattacharya_2023 ; Chakrabarti:2025hsb ; NSSrivatsa:2023pby :
$\mathcal{L}=\begin{pmatrix}a_{0}&b_{0}&0&\cdots\\ c_{0}&a_{1}&b_{1}&\ddots\\ 0&c_{1}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$,
where $a_{n},b_{n},c_{n}\in\mathbb{C}$. The goal is to transform $\mathcal{L}$ such that the off-diagonals become equal, specifically $\sqrt{b_{n}c_{n}}$, to facilitate moment-based Krylov complexity calculations gaaf2016infinitebilanczosmethodnonlinear .
Define a diagonal matrix $D=\text{diag}(d_{0},d_{1},\dots)$ with $d_{0}=1$ and ratios $d_{k+1}/d_{k}=\sqrt{c_{k}/b_{k}}$, choosing the principal branch of the square root (argument in $(-\pi/2,\pi/2]$) for consistency. The transformed Lindbladian is $\mathcal{L}^{\prime}=D^{-1}\mathcal{L}D$, which preserves eigenvalues and has the form Yamamoto1968 :
$\mathcal{L}^{\prime}=\begin{pmatrix}a_{0}&\sqrt{b_{0}c_{0}}&0&\cdots\\ \sqrt{b_{0}c_{0}}&a_{1}&\sqrt{b_{1}c_{1}}&\ddots\\ 0&\sqrt{b_{1}c_{1}}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$.
For complex values of $\sqrt{b_{n}c_{n}}$ gaaf2016infinitebilanczosmethodnonlinear , these terms can be written in modulus–argument form. In this representation, the equal off-diagonal elements share constant phases, which are absorbed by the transformation matrix $D$. This redefines the basis up to an overall phase, so that only the moduli of the coefficients need to be considered. Note that, in dissipative systems, such as the SYK model with Lindblad operators Bhattacharjee:2023uwx , $b_{n}=c_{n}$ may be real and positive, simplifying to $\sqrt{b_{n}c_{n}}=b_{n}$. For general cases, the transformation symmetrizes the off-diagonals, aiding moment computations for Krylov complexity while accounting for dissipative effects, where the imaginary parts of $a_{n}$ typically suppress complexity growth.
In this transformed basis the Lanczos coefficients $\{b_{n}\}$ are replaced by $\sqrt{b_{n}c_{n}}$ and are related to the moments $\{\mu_{n}\}$ as NSSrivatsa:2023pby :
$b_{n}c_{n}$ $=\frac{\tau_{n+1}(t)\tau_{n-1}(t)}{\tau_{n}^{2}(t)},\quad n\geq 0.$ (12)
In the following, we first consider a damped harmonic oscillator, followed by the well-known Caldeira–Leggett model as our representative open quantum systems. We then apply the moments method to compute and study the evolution of Krylov complexity.
3 Damped Harmonic Oscillator
The first model we consider is the well-known damped harmonic oscillator. Its Hamiltonian describes a harmonic oscillator coupled to an electromagnetic field, serving as a representative example of damping:
$H=\omega_{1}a^{\dagger}a+\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}+\sum_{k}\gamma_{k}(a^{\dagger}b_{k}+ab_{k}^{\dagger}).$ (13)
Unlike closed quantum systems, open quantum systems do not evolve unitarily. Instead, their dynamics are governed by equations of motion for the density matrix, which generalize classical Markovian processes. This formalism leads to the Lindblad master equation:
$\frac{d\rho_{s}(t)}{dt}=-i[H,\rho_{s}(t)]+\sum_{i}\left(L_{i}^{\dagger}\rho_{s}(t)L_{i}-\frac{1}{2}L_{i}^{\dagger}L_{i}\rho_{s}(t)-\frac{1}{2}\rho_{s}(t)L_{i}^{\dagger}L_{i}\right).$ (14)
This equation assumes a Markovian bath, where bath correlations are proportional to delta functions in time, implying a memoryless environment where information lost to the bath cannot return to the system. This is central to dissipative dynamics. The adjoint master equation, describing operator evolution, is:
$\frac{dA}{dt}=i[H,A]+\sum_{i}\left(L_{i}^{\dagger}AL_{i}-\frac{1}{2}L_{i}^{\dagger}L_{i}A-\frac{1}{2}AL_{i}^{\dagger}L_{i}\right)\,.$ (15)
The first term generates unitary evolution, while the remaining terms account for transitions due to interactions with the bath. The operators $L_{i}$ are jump operators, with $L_{i}^{\dagger}AL_{i}$ driving quantum transitions and the terms $\frac{1}{2}L_{i}^{\dagger}L_{i}A$ and $\frac{1}{2}AL_{i}^{\dagger}L_{i}$ correcting for cases where no jump occurs Preskill1998 .
Using the framework mentioned in Sec. (2), we study the Krylov complexity in this dissipative system. The jump operators are chosen to maintain the Markovian approximation. In open systems, information and energy flow bidirectionally between the system and bath, leading to non-Markovian fluctuations. A good Markovian approximation requires sufficient separation between the correlation time of fluctuations and the system’s evolution timescale. We model the system-bath interaction as:
$H_{I}=\sum_{k}\gamma_{k}(ab_{k}^{\dagger}+a^{\dagger}b_{k}).$ (16)
Assuming the environment is initially at zero temperature, the system can emit energy without absorbing any from the bath. We define a single Lindblad operator for photon emission, proportional to the annihilation operator $a$:
$L=\sqrt{\gamma}a,$ (17)
where $\gamma$ is the decay rate from the first excited state to the ground state. The master equation for a system operator $O(t)$ becomes:
$\dot{O}(t)=i[H_{0},O(t)]+\gamma\left(a^{\dagger}O(t)a-\frac{1}{2}a^{\dagger}aO(t)-\frac{1}{2}O(t)a^{\dagger}a\right)$ (18)
where $H_{0}=\omega_{1}a^{\dagger}a$ is the standard harmonic oscillator Hamiltonian. For the Caldeira-Leggett model, we will consider a more generic quadratic form for the operator $O=\frac{1}{2\omega}(a^{\dagger}a^{\dagger}+aa+a^{\dagger}a+aa^{\dagger})$, which then evolves as:
$O(t)=\frac{1}{2\omega}\left(f_{1}(t)a^{\dagger}a^{\dagger}+f_{2}(t)aa+f_{3}(t)a^{\dagger}a+f_{4}(t)aa^{\dagger}\right),$ (19)
where $\{f_{1},f_{2},f_{3},f_{4}\}\in\mathbb{C}$. Solving the differential equation for the individual annihilation and creation operators $a(t)$ and $a^{{\dagger}}(t)$, we obtain:
$a^{{\dagger}}(t)$ $=e^{i\omega t-\frac{\gamma}{2}t}a^{{\dagger}}\,$, (20)
$a(t)$ $=e^{-i\omega t-\frac{\gamma}{2}t}a\,.$ (21)
Combining these yields the following:
$O(t)=\frac{1}{2\omega}(e^{-i2\omega t-\gamma t}aa+e^{i2\omega t-\gamma t}a^{\dagger}a^{\dagger}+e^{-\gamma t}a^{\dagger}a+e^{-\gamma t}aa^{\dagger})\,.$ (22)
This result is consistent with the harmonic oscillator, with additional exponential decay terms reflecting information loss to the environment. The thermal two-point correlation function is:
$\langle Q(t)Q(0)\rangle_{\beta\rightarrow\infty}=2e^{-i2\omega t-\gamma t}-e^{-\gamma t}.$ (23)
Using the moments method from Sec. (2), we compute the probability amplitudes $(|\phi_{n}|^{2})$ and the corresponding Krylov complexity for the under and overdamped cases, which are displayed in Fig. (1) and (2) respectively for different values of $\omega$:
Figure 1: Time evolution of Krylov complexity (K-complexity) for under-damped ($\omega=1\,,\gamma=0.1$) and over-damped systems ($\omega=1\,,\gamma=0.9$).
Figure 2: Time evolution of probability amplitudes for under-damped ($\omega=1\,,\gamma=0.1$) and over-damped ($\omega=1\,,\gamma=0.9$) systems.
From the Fig. (1), we observe that at early times Krylov complexity exhibits rapid growth, followed by a decay toward a low saturation value. This behavior contrasts with the harmonic oscillator results presented in the Fig. (8) of the Appendix (A). Unlike the harmonic oscillator, which is a closed system, the damped oscillator loses information to the environment without any backflow. For the harmonic oscillator, the complexity displays periodic oscillations between zero and one, reflecting the evolution of the position operator into the momentum operator and vice versa, as dictated by the canonical commutation relations. These periodic oscillations are absent in the damped case.
In line with the conjecture made in Liu_2023 , dissipation leads to a lower overall complexity and a reduced saturation value. We therefore identify the oscillatory decay of Krylov complexity to a low saturation value as a hallmark of dissipation. Furthermore, the maximum complexity reached is significantly smaller than in the harmonic oscillator, as can be seen from the Fig. (1) and (8). This suggests that, in the overdamped regime, the operator explores only a limited portion of the system space before dissipating into the bath, while in the underdamped regime, it explores more of the system space (compared to the overdamped case, as evidenced by its higher saturation value) prior to this loss. Nevertheless, in both cases the spread into the bath occurs on the same timescale, consistent with the observed saturation. This interpretation is further supported by the probability amplitude plots as shown in Fig. (2): in the underdamped case, the slower decay indicates a less rapid loss of information to the bath.
4 Caldeira-Leggett Model
The Caldeira-Leggett model is a system-bath model designed to study quantum Brownian motion CALDEIRA1983374 ; Caldeira:1982iu ; PhysRevLett.46.211 . It describes a Brownian particle of mass $m$ coupled via coordinate-coordinate coupling to a bath of non-interacting harmonic oscillators:
$H=\frac{p^{2}}{2m}+\frac{m\omega^{2}x^{2}}{2}+\sum_{n}\left(\frac{p_{n}^{2}}{2m_{n}}+\frac{m_{n}\omega_{n}^{2}x_{n}^{2}}{2}\right)+x\sum_{n}\kappa_{n}x_{n}.$ (24)
In the Schrödinger picture, the dynamics of the reduced system are described by the Caldeira-Leggett master equation BRE02 :
$\frac{d}{dt}\rho_{s}(t)=-\frac{i}{\hbar}[H_{s},\rho_{s}(t)]-\frac{i\gamma}{\hbar}[x,\{p,\rho_{s}(t)\}]-\frac{2m\gamma k_{B}T}{\hbar^{2}}[x,[x,\rho_{s}(t)]],$ (25)
where $H_{s}=\frac{p^{2}}{2m}+\frac{m\omega^{2}x^{2}}{2}$. The first term describes coherent dynamics, the second term, proportional to the relaxation rate $\gamma$, drives dissipation by damping momentum expectation values, and the third term, proportional to temperature $T$, induces decoherence through thermal fluctuations Schlosshauer_2019 ; BRE02 . The Caldeira-Leggett master equation is not directly in Lindblad form, as its dissipator is not positive definite. In the high-temperature limit, this can be remedied by adding a minimally invasive term $\left(-\frac{\gamma}{8mk_{B}T}[p,[p,\rho_{s}(t)]]\right)$ Ramazanoglu_2009 ; BRE02 ; Schlosshauer_2019 . The Lindbladian form facilitates analysis of operator dynamics in the adjoint form:
$\frac{d}{dt}A(t)=i\left[H_{s}+\frac{\gamma}{2}(xp+px),A\right]+\gamma\left(L^{\dagger}AL-\frac{1}{2}L^{\dagger}LA-\frac{1}{2}AL^{\dagger}L\right),$ (26)
with the Lindblad jump operator defined as
$L=Ax+iBp$ (27)
where $A=\sqrt{\frac{4mk_{B}T}{\hbar^{2}}}$ and $B=\sqrt{\frac{1}{4mk_{B}T}}$. The approximate time evolution for an operator $O(t)$ is:
$\frac{d}{dt}O(t)$ $\propto-\omega^{2}(xp+px)-2\gamma p^{2}+\gamma(A^{2}-2ABp^{2}),$ (28)
which in terms of creation and annihilation basis can be written as
$\frac{d}{dt}O$ $=-\omega^{2}(xp+px)+(-2\gamma-2\gamma AB)p^{2}+\gamma A^{2}(aa^{\dagger}-a^{\dagger}a).$ (29)
Here the proportionality factor is taken to be unity. Factorizing the time dependence of the creation and annihilation operators as $a(t)=f^{*}(t)a$ and $a^{\dagger}(t)=f(t)a^{\dagger}$, and solving the differential equation we obtain the following time evolved operator
$O(t)$ $=-\frac{\omega}{2}\Big(e^{(2i\omega-2\gamma-2\gamma AB)t}aa-e^{(\gamma A^{2}-2\gamma-2\gamma AB)t}aa^{\dagger}$
$-e^{(-\gamma A^{2}-2\gamma-2\gamma AB)t}a^{\dagger}a+e^{(-2i\omega-2\gamma-2\gamma AB)t}a^{\dagger}a^{\dagger}\Big).$ (30)
Assuming the system is in thermal equilibrium, the density matrix $\rho_{\beta}$ is taken to be a thermal Gibbs state of the form $\rho_{\beta}=\frac{e^{-\beta H}}{Z}$. The auto-correlation function with respect to the thermal average is then given by:
$\langle O(t)O\rangle_{\beta}$ $=\frac{\mathrm{Tr}\{\rho_{\beta}O(t)O\}}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\,$, (31)
$=e^{(2i\omega-2\gamma-2\gamma AB)t}(C_{3}+3C_{2}+2C_{1})+e^{(\gamma A^{2}-2\gamma-2\gamma AB)t}(2C_{3}+3C_{2}+C_{1})$
$\quad+e^{(-\gamma A^{2}-2\gamma-2\gamma AB)t}(2C_{3}+C_{2})+e^{(-2i\omega-2\gamma-2\gamma AB)t}(C_{3}-C_{2})$ (32)
where the coefficients $C_{1},C_{2}$ and $C_{3}$ are given as,
$C_{1}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}e^{-\beta(n+1/2)\omega}=\frac{(-1+e^{\beta\omega})^{2}}{3(e^{\beta\omega}+1)^{2}}\,$, (33)
$C_{2}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}ne^{-\beta(n+1/2)\omega}=\frac{-1+e^{\beta\omega}}{3(e^{\beta\omega}+1)^{2}}\,$, (34)
$C_{3}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}n^{2}e^{-\beta(n+1/2)\omega}=\frac{1}{3+3e^{\beta\omega}}\,.$ (35)
These coefficients and the two-point function are evaluated at various temperatures, consistent with the high-temperature limit of the Caldeira–Leggett master equation.
To investigate the role of dissipation and correlations, we begin in the following subsections with an analysis of the decoherence function, followed by a study of Krylov complexity. Our aim is to determine whether Krylov complexity is sensitive to decoherence in a manner similar to circuit complexity Bhattacharyya_2024 ; Bhattacharyya_2023 .
4.1 Decoherence Function
Interactions between an open quantum system and its environment create correlations between states of the system and the environment. In specific basis these correlations/coherence in the density matrix are strongly decaying, a process tracked by the decoherence function, which describes the behavior of the off-diagonal elements of the reduced density matrix. Consider, as the starting point of our discussion, the Hamiltonian:
$H=H_{S}+H_{B}+H_{I}=H_{0}+H_{I},$ (36)
The interaction Hamiltonian takes the following form,
$H_{I}=\sum_{n}|n\rangle\langle n|\otimes B_{n}\equiv\sum_{n}A_{n}\otimes B_{n}.$ (37)
This interaction Hamiltonian singles out a specific set of basis vectors of the system and $B_{n}$ denote reservoir operators, the choice of $A_{n}=|n\rangle\langle n|$ is such that $A_{n}$ are conserved quantities, i.e. that
$H_{0}+H_{I},A_{n}]=[H_{0},A_{n}]=0\,.$ (38)
By this condition, it is clear that the Krylov basis is not a suitable basis for deriving the decoherence function. As a requirement of the Lanczos algorithm, the commutation relations between the Hamiltonian and the operator are non-trivial; otherwise, (3) would terminate instantly. Then, from (38), we can see that in the interaction picture, the time evolution of the interaction Hamiltonian passes entirely onto the bath operators.
$H_{I}(t)=e^{iH_{0}t}H_{I}e^{-iH_{0}t}=\sum_{n}|n\rangle\langle n|\otimes B_{n}(t)\,.$ (39)
As a result of the commutation relation the basis states $|n\rangle$ are not effected by the coupled dynamics. This gives us the following system bath state:
$|\Psi(t)\rangle=\sum_{n}c_{n}|n\rangle\otimes|\phi_{n}(t)\rangle\,.$ (40)
It represents correlations between states describing the system and the corresponding states of the reservoir. Due to the interaction, the reservoir carries information on the system state. These coherences/correlations are still present in the system’s reduced density matrix:
$\rho_{s}(t)=\mathrm{Tr}_{B}{|\Psi(t)\rangle\langle\Psi(t)|}=\sum_{n}\sum_{m}c_{n}c^{*}_{m}|n\rangle\langle m|\langle\phi_{m}(t)|\phi_{n}(t)\rangle\,.$ (41)
It is clear that the diagonal elements of $\rho_{s}(t)$ are constant in time while the off-diagonals vary in time. The time dependence of $\rho_{s}(t)$ is given by the overlap between the corresponding reservoir states:
[…]
The quantity $\Gamma_{nm}(t)$ characterizes the behaviour of the off diagonal terms in the reduced density matrix and is known as the decoherence function. For the Caldeira-Leggett model, the decoherence function is BRE02 :
$\Gamma(t)$ $=-\frac{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}{8\sigma_{0}^{2}}\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}},$ (43)
$\tilde{\Gamma}(t)$ $=\frac{8\sigma_{0}^{2}\Gamma(t)}{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}=\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}},$ (44)
where $I$, $G_{1}$, and $G_{2}$ are the inhomogeneous and homogeneous solutions of the operator $x(t)$ evolution, and $\Delta x$ and $\Delta p$ are the initial separations of superposed Gaussian wave packets in position and momentum space, respectively.
Figure 3: Time evolution of decoherence function $\tilde{\Gamma}$ for $\omega=1$ and $\gamma=0.1.$ .
From Fig. (3), it is clear that the decoherence rate saturates, as described in Schlosshauer_2019 . This occurs when $x-x^{\prime}$ reaches the maximum coherence length of the oscillator environment. This length is the characteristic scale over which the phase coherences of the environment, and hence the bath correlations, remain significant. The saturation of the decoherence function implies that the off-diagonal elements of the density matrix vanish. However, this feature is basis-dependent, and as noted earlier in the section, the Krylov basis is not the appropriate one for capturing it. We therefore suspect that Krylov complexity may be insensitive to decoherence, a possibility we will explore in the next subsection.
4.2 Krylov Complexity
As outlined in Sec. (2), we use the moments method to compute the Krylov complexity for the Caldeira–Leggett model. To investigate the roles of dissipation and correlation, we examine the time evolution of the Krylov complexity, shown in Fig. (4).
Figure 4: Time evolution of Krylov complexity for the Caldeira-Leggett master equation with $\omega=1$ and $\gamma=1/10\,.$ On the right, the late time complexity for the same values of $\gamma$ and $\omega$ is displayed.
Figure 5: Time evolution of Krylov complexity for the Caldeira-Leggett master equation with $\omega=1/10$ and $\gamma=1/20$. On the right, the saturation of the complexity for the same values of $\gamma$ and $\omega$ is displayed.
From Fig. (4), we observe two key features of the Krylov complexity in the Caldeira-Leggett model:
  • The initial oscillatory behavior is accompanied by a secondary frequency mode—‘anti-oscillations,’ which appear at the peaks of each oscillation, as shown in Fig. (4) and Fig. (5). This feature causes the Krylov complexity to behave differently from the results in Sec. (3) with dissipation. In the damped oscillator case studied in Sec. (3), the contribution from decoherence was suppressed, unlike in the Caldeira–Leggett model. As a result, the Krylov complexity here exhibits a much slower decay compared to Sec. (3).
  • The complexity saturates, a well-known result for open quantum systems Bhattacharya_2023 ; Liu_2023 . This saturation reflects the interplay between the decoherence and dissipative terms in the master equation.
For a given set of parameters, we find that the complexity saturates at a value close to $0.2$, within an interval of $0.1$, as illustrated in Fig. (5). This saturation occurs at approximately $t=93$. The saturation is driven by the interplay of dissipation and decoherence. To confirm this, we further investigate the model in the following analysis.
4.3 Decomposing the Master Equation into Individual Terms
Analzing the master equation it is clear that under particular operator choices and assumptions we can isolate the contributions from individual terms in the master equation and study how these terms contribute to Krylov complexity. Therefore in the following sections we will consider particular choices that allows us to investigate the individual contributions of decoherence and dissipation.
Decoupling Dissipation from the Master Equation
To separate the effects of dissipation on Krylov complexity from the master equation, we turn off the term proportional to the relaxation rate $\gamma$, thereby considering the Lindblad double-commutator form of the master equation Schlosshauer_2019 ; BRE02  .
$\frac{d}{dt}\rho_{s}(t)=-\frac{i}{\hbar}[H_{s},\rho_{s}(t)]-\frac{2m\gamma k_{B}T}{\hbar^{2}}[x,[x,\rho_{s}(t)]].$ (45)
This yields a master equation that accounts only for the coherent system dynamics and the environment-induced decoherence, as discussed in Ferialdi_2017 ; BRE02 ; Schlosshauer_2019 .
Figure 6: Left: Time evolution of Krylov complexity (K-complexity) for $\omega=1$ and $\gamma=1/100$. Right: Time evolution of Krylov complexity for the same parameters, with the decoherence function $\tilde{\Gamma}(t)$ overlaid in orange.
One significant difference we immediately notice from Fig. (6), is the absence of saturation. Note that we read off the saturation timescale from the decoherence function, and at the same timescale, the Krylov complexity, computed for identical parameters, shows diminishing oscillations. However, since the pattern remains periodic, it cannot be taken as a definitive signature of the onset of decoherence. The varying saturation timescales in Fig. (4) and Fig. (5) arise from the competition between the dissipative and decoherence terms. In the case where the damping coefficient $\gamma$ is large compared to $\omega$ (Fig. (5)), the saturation timescale is significantly shorter. The decoherence term generates the secondary oscillation (the “anti-oscillations” near each of the peaks) as observed in Fig. (5). We further checked that as we increased the temperature, the period of these oscillations in the Krylov complexity became shorter, appearing as “noise”.
Our analysis of complexity shows no clear evidence for the decay of system–bath correlations. Note that in the Krylov basis, the Lindbladian takes a tridiagonal structure characterized by the Lanczos coefficients.
$\frac{d}{dt}\rho_{s}(t)=\underbrace{\begin{pmatrix}a_{0}&\sqrt{b_{0}c_{0}}&0&\cdots\\ \sqrt{b_{0}c_{0}}&a_{1}&\sqrt{b_{1}c_{1}}&\ddots\\ 0&\sqrt{b_{1}c_{1}}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}.}_{\mathcal{L}}\rho_{s}(t),$ (46)
For $\rho_{s}(t)$ in the form given by (41), it is evident that at least one of the diagonal elements of the reduced density matrix is time dependent, while at least one of the off-diagonal elements contains a term without time dependence. Consequently, there is no clear separation of the correlations carried by the off-diagonal terms, and the Krylov complexity cannot be reliably used to indicate the decay of coherences.
Finally, the inclusion of the dissipative term determines how rapidly the complexity decays or grows. As shown in Fig. (6), for the same set of parameters, the complexity exhibits significantly different global features compared to Fig. (5).
Decoupling Decoherence
To eliminate the effects of decoherence, we select a suitable operator that removes the temperature-dependent contribution. In particular, by choosing $O=x^{2}$, as in Sec. (3), the decoherence term in (25) is suppressed. The two-point function then becomes:
$\langle O(t)O\rangle_{\beta}$ $=\left(e^{(4i\omega-2\gamma-2\gamma AB)t}(C_{3}+3C_{2}+2C_{1})+e^{(2\omega\gamma B^{2}-2\gamma-2\gamma AB)t}(2C_{3}+3C_{2}+C_{1})\right$.
$\quad\left.+e^{(-2\omega\gamma B^{2}-2\gamma-2\gamma AB)t}(2C_{3}+C_{2})+e^{(-4i\omega-2\gamma-2\gamma AB)t}(C_{3}-C_{2})\right).$ (47)
This operator choice underscores the role of the Lindblad jump operator coefficient $B$, which is proportional to the momentum, in governing the system’s dissipative dynamics. In the high-temperature limit, the $B^{2}$ terms are suppressed, recovering the dissipative behavior discussed in Sec. (3), characterized by sharp complexity growth followed by decaying oscillations toward a small saturation value. The decoherence term, however, introduces noise that manifests as oscillations around the saturation point.
Figure 7: Time evolution of Krylov complexity and probability amplitudes for the Caldeira-Leggett master equation with operator $O=x^{2}$, $\omega=1$, and $\gamma=1/10$.
5 Discussion
In this work, we investigate Krylov complexity in open quantum systems, focusing on the damped harmonic oscillator and the Caldeira–Leggett model. Our results show that Krylov complexity effectively captures dissipative and decoherence effects, revealing distinct behaviors in these systems. In the Caldeira–Leggett model, the late-time Krylov complexity exhibits saturation, indicative of a fully mixed state driven by dissipation and decoherence, but it does not clearly reflect the onset of decoherence. In contrast, for the damped harmonic oscillator, where decoherence is suppressed, Krylov complexity highlights pronounced dissipative effects, characterized by a sharp initial increase followed by oscillatory decay. These findings underscore the potential of Krylov complexity as a diagnostic tool for probing open quantum systems and open avenues for further exploration in quantum information theory and quantum Brownian motion.
Our analysis of the Caldeira-Leggett master equation reveals significant differences in the contributions of its individual terms, which account for dissipation and decoherence. As a preliminary model, we employed the damped harmonic oscillator master equation to isolate the features of dissipation in Krylov complexity. In this case, Krylov complexity exhibits a rapid initial rise followed by an oscillatory decay, reflecting the dissipative dynamics. In the Caldeira–Leggett model, which contains distinct terms for dissipation and decoherence, we found that with appropriate choices for the operator and in the high-temperature limit, these terms can be independently suppressed. When the decoherence term is turned off, the results mirror those of the damped harmonic oscillator, showing a sharp increase and oscillatory decay driven by dissipation. Conversely, when the dissipative term is suppressed, the dynamics align with the coherent behavior described in Appendix (A). In this scenario, the decoherence term in the Caldeira-Leggett master equation primarily drives operator diffusion rather than dissipation, leading to oscillatory behavior with periodic “bottlenecks” caused by a second oscillatory frequency introduced by decoherence. Increasing the temperature in the decoherence term amplifies the frequency of these oscillations, manifesting as “noise” in the Krylov complexity, characterized by rapid oscillations and a slow decay due to the interplay of the second frequency mode.
We compare our Krylov complexity results for the Caldeira–Leggett model with those of Bhattacharyya_2023 , where circuit complexity based on Nielsen’s geometric approach was used to compute the complexity of purification for the reduced density matrix in a system coupled to a free scalar field bath. Both Krylov and circuit complexity exhibit saturation, signaling a fully mixed state due to decoherence. In Bhattacharyya_2023 , it was noted that circuit complexity saturates on a timescale similar to that of linear entropy and entanglement negativity. By contrast, our Krylov complexity results saturate at a later stage, once decoherence becomes dominant, compared with the onset indicated by the decoherence function. This suggests that circuit complexity is more sensitive to detecting the onset of mixedness, while Krylov complexity highlights operator growth as decoherence progresses.
This relative insensitivity of Krylov complexity to the early stages of decoherence can be traced to the choice of basis. In particular, the behavior of the off-diagonal terms in the density matrix strongly depends on the so-called preferred basis Schlosshauer_2019 ; BRE02 . The Krylov basis required for the Krylov complexity does not satisfy the conditions of the preferred basis, which may explain why our results for the Caldeira–Leggett master equation do not clearly reproduce the decoherence signatures observed in prior studies. Future work could explore how operator growth defined in the preferred basis might better capture decoherence signals, thereby enhancing the utility of operator complexity in open quantum systems.
In parallel with this line of inquiry, related extensions are also underway. For instance, in Bhattacharyya_2024 , the Complexity of Purification was analyzed for the Caldeira–Leggett model in curved space, revealing additional distinctive features. It would be interesting to examine whether similar behavior also appears in the context of Krylov operator complexity, and we are currently investigating this question. Another promising direction is to study recoherence phenomena in these curved-space models Bhattacharyya:2025cxv .
Acknowledgments
The authors would like to thank HJR Van Zyl and Arpon Paul for discussions. SSH is supported in part by the National Institute for Theoretical and Computational Sciences of South Africa (NITheCS). AB is supported by the Core Research Grant (CRG/2023/ 001120) by the Department of Science and Technology (DST) and Anusandhan National Research Foundation (ANRF), India. MSG is supported by the National Research Foundation postgraduate bursary.
Appendix A Krylov Complexity for the Harmonic Oscillator
Recall that from Eq. (2) the operator evolution is given as:
$O(t)$ $=O+(it)[H,O]+\frac{(it)^{2}}{2!}[H,[H,O]]+\frac{(it)^{3}}{3!}[H,[H,[H,O]]]+\ldots$ (48)
We begin with the operator $O=\alpha(a^{{\dagger}}+a)$ and Hamiltonian $H=\omega(a^{\dagger}a+\frac{1}{2})$. Then the first few commutators in Eq. (48) is calculated as:
$[H,O]$ $=\omega\alpha(a^{{\dagger}}-a)\,$, (49)
$[H,[H,O]]$ $=\omega^{2}\alpha(a^{{\dagger}}+a)$ (50)
$,,$ (51)
$[H,[H,[H,O]]]$ $=\omega^{3}\alpha(a^{{\dagger}}-a)$ (52)
from this we can see that there is clearly a pattern. Then inserting this back into Eq. (48) we get:
$O(t)$ $=\alpha(a^{{\dagger}}+a)+(it\omega)\alpha(a^{{\dagger}}-a)+\frac{(it\omega)^{2}}{2!}\alpha(a^{{\dagger}}+a)+\frac{(it\omega)^{3}}{3!}\alpha(a^{{\dagger}}-a)+\ldots\,$, (53)
$=\left[\alpha(a^{{\dagger}}+a)\left(1+\frac{(it\omega)^{2}}{2!}+\ldots\right)\right]+\left[\alpha(a^{{\dagger}}-a)\left(it\omega+\frac{(it\omega)^{3}}{2!}+\ldots\right)\right]\,$, (54)
$=\alpha\cos(\omega t)(a^{{\dagger}}+a)+i\alpha\sin(\omega t)(a^{{\dagger}}-a)$ (55)
$=\cos(\omega t)x+\frac{1}{m\omega}\sin(\omega t)p\,$, (56)
$=\alpha(e^{i\omega t}a^{{\dagger}}+e^{-i\omega t}a)\,.$ (57)
Running this through the procedure of the moments method mentioned in Section (2), as well as the Lanczos algorithm we find the following result for the Krylov complexity which we show in Fig. (8).
Figure 8: Time evolution of Krylov complexity for Harmonic oscillator with $\omega=1$ and $m=1$.

### v2

@@
 (5), we summarize our findings and discuss their broader implications for future research.
-In the Appendix (A), we have provided some additional details.
+In the Appendix, we have provided some additional details.
 2 Krylov Complexity
@@
 […]
-Krylov Complexity for Open Quantum System
-For open quantum systems governed by a Lindbladian $\mathcal{L}$, which describes non-unitary evolution due to environmental coupling, the standard Lanczos algorithm for Hermitian operators is insufficient.
-Instead, the bi-Lanczos algorithm is required to handle the non-Hermitian nature of the Lindbladian Bhattacharya_2023 ; Chakrabarti:2025hsb .
-This algorithm constructs a bi-orthogonal Krylov basis, yielding a tridiagonal matrix representation with complex diagonal elements $a_{n}$ and off-diagonal elements $b_{n}$ (superdiagonal) and $c_{n}$ (subdiagonal).
-These off-diagonals are generally complex and unequal ($b_{n}\neq c_{n}^{*}$), reflecting the dissipative dynamics.
-In moment-based methods for computing Krylov complexity, the product $b_{n}c_{n}$ often appears, but its complex nature can complicate calculations, especially when taking the square root for complexity metrics.
-A similarity transformation can symmetrize the off-diagonals to $\sqrt{b_{n}c_{n}}$, simplifying moment computations while preserving the spectrum.
-The Lindbladian’s tridiagonal form in the bi-Lanczos basis is Bhattacharya_2023 ; Chakrabarti:2025hsb ; NSSrivatsa:2023pby :
-$\mathcal{L}=\begin{pmatrix}a_{0}&b_{0}&0&\cdots\\ c_{0}&a_{1}&b_{1}&\ddots\\ 0&c_{1}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$,
+Figure 1: Schematic diagram of the spreading of the operator, as it evolves, in operator space.
+Krylov Complexity for Open Quantum Systems
+For open quantum systems governed by a Lindbladian $\mathcal{L}$ (a generalized Liouvillian), which describes non-unitary evolution due to environmental coupling, the standard Lanczos algorithm for Hermitian operators is insufficient.
+Instead, there are a variety of ways in which one can treat such quantum systems using Krylov methods.
+Among them are the Arnoldi algorithm, Closed Krylov basis and the bi-Lanczos algorithm.
+The closed Krylov basis Liu_2023 is an identical treatment to the generalization of the closed system dynamics for a non-Hermitian initial operator, as was discussed in the previous sub-section.
+Both the Arnoldi and bi-Lanczos algorithms attempt to generalize the Lanczos algorithm for open systems.
+The Arnoldi algorithm does this by constructing an orthonormal basis; however, it turns out that in this basis, the Lindbladian is not tri-diagonal, instead it is in an upper Hessenberg form.
+On the other hand the bi-Lanczos algorithm aims to preserve the tri-diagonal structure obtained in the Lanczos algorithm.
+This is done by evolving two initial vectors $|p_{0}\rangle$ and $|q_{0}\rangle$ by the adjoint Lindbladian $\mathcal{L^{\dagger}}$ and the Lindbladian $\mathcal{L}\,$, respectively.
+As such, it creates a bi-orthonormal basis instead of an orthonormal one,
+$\mathrm{Kry}^{j}(\mathcal{L^{\dagger}},|p_{0}\rangle)$ $=\{|p_{0}\rangle,\mathcal{L^{\dagger}}|p_{0}\rangle,(\mathcal{L^{\dagger}})^{2}|p_{0}\rangle,\ldots\},$ (12)
+$\mathrm{Kry}^{j}(\mathcal{L},|q_{0}\rangle)$ $=\{|q_{0}\rangle,\mathcal{L}|q_{0}\rangle,(\mathcal{L})^{2}|q_{0}\rangle,\ldots\},$ (13)
+with the bi-orthonormality condition, $\langle q_{m}|p_{n}\rangle=\delta_{mn}$ Bhattacharya_2023; Chakrabarti:2025hsb.
+Unlike the Arnoldi algorithm, which leaves the Linbdladian in an upper Hessenberg form, the bi-Lanczos algorithm, as mentioned earlier, preserves the tri-diagonal structure with complex diagonal elements $a_{n}$ and off-diagonal elements $b_{n}$ (super-diagonal) and $c_{n}$ (sub-diagonal).
+Generated by equivalent statements of (12), by the action of $L$, $L^{{\dagger}}$ on the vectors.
+$\mathcal{L^{{\dagger}}}|p_{j}\rangle$ $=b_{j}|p_{j-1}\rangle+a_{j}|p_{j}\rangle+c_{j+1}|p_{j+1}\rangle\,$, (14)
+$\mathcal{L}|q_{j}\rangle$ $=c^{*}_{j}|q_{j-1}\rangle+a^{*}_{j}|q_{j}\rangle+b^{*}_{j+1}|q_{j+1}\rangle\,.$ (15)
+Using this representation, the matrix elements of the Lindbladian can be written as $\langle q_{i}|\mathcal{L^{{\dagger}}}|p_{j}\rangle$, allowing it to be expressed as:
+$\mathcal{L^{{\dagger}}}=\begin{pmatrix}a_{0}&b_{0}&0&\cdots\\ c_{0}&a_{1}&b_{1}&\ddots\\ 0&c_{1}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$ (16)
 where $a_{n},b_{n},c_{n}\in\mathbb{C}$.
-The goal is to transform $\mathcal{L}$ such that the off-diagonals become equal, specifically $\sqrt{b_{n}c_{n}}$, to facilitate moment-based Krylov complexity calculations gaaf2016infinitebilanczosmethodnonlinear .
-Define a diagonal matrix $D=\text{diag}(d_{0},d_{1},\dots)$ with $d_{0}=1$ and ratios $d_{k+1}/d_{k}=\sqrt{c_{k}/b_{k}}$, choosing the principal branch of the square root (argument in $(-\pi/2,\pi/2]$) for consistency.
+Using a similarity transformation, the off-diagonal entries can be written as a generalized Lanczos coefficient, $\tilde{b}_{n}=\sqrt{b_{n}c_{n}}$.
+Define a diagonal matrix $D=\text{diag}(d_{0},d_{1},\dots)$ with $d_{0}=1$ and ratios $d_{k}=\sqrt{c_{k}/b_{k}}$.
 The transformed Lindbladian is $\mathcal{L}^{\prime}=D^{-1}\mathcal{L}D$, which preserves eigenvalues and has the form Yamamoto1968:
 $\mathcal{L}^{\prime}=\begin{pmatrix}a_{0}&\sqrt{b_{0}c_{0}}&0&\cdots\\ \sqrt{b_{0}c_{0}}&a_{1}&\sqrt{b_{1}c_{1}}&\ddots\\ 0&\sqrt{b_{1}c_{1}}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$.
+(17)
+The benefit of this similarity transformation is that it allows us to generalize the moments method for closed systems to open systems.
+In general, the Lanczos coefficients $b_{n}$ are taken to be positive real numbers, with the diagonal coefficients generally being complex.
 For complex values of $\sqrt{b_{n}c_{n}}$ gaaf2016infinitebilanczosmethodnonlinear, these terms can be written in modulus–argument form.
-In this representation, the equal off-diagonal elements share constant phases, which are absorbed by the transformation matrix $D$.
+In this representation, the equal off-diagonal elements share constant phases, which are absorbed by the transformation matrix $D_{1}$.
+$D_{1}=\mathrm{diag}\left(1,\exp\left\{i\phi_{1}\right\},\exp\left\{i(\phi_{1}+\phi_{2}\right\},\ldots,\exp\left\{i\sum_{n}\phi_{n}\right\}\right)$ (18)
+Such that,
+$\mathcal{L}^{\prime}=D^{-1}_{1}\mathcal{L}D_{1}\,.$ (19)
 This redefines the basis up to an overall phase, so that only the moduli of the coefficients need to be considered.
+This is possible as the diagonal coefficients $a_{n}$ are the ones that ensure the non-Hermitian nature of the Lindbladian Nandy_2025.
+In the absence of the complex diagonal coefficients, the Lindbladian/Liouvillian should be hermitian with the off-diagonal coefficients being, in the most general case, complex conjugate pairs caputa2021geometrykrylovcomplexity; Dymarsky_2021.
 Note that, in dissipative systems, such as the SYK model with Lindblad operators Bhattacharjee:2023uwx, $b_{n}=c_{n}$ may be real and positive, simplifying to $\sqrt{b_{n}c_{n}}=b_{n}$.
-For general cases, the transformation symmetrizes the off-diagonals, aiding moment computations for Krylov complexity while accounting for dissipative effects, where the imaginary parts of $a_{n}$ typically suppress complexity growth.
-In this transformed basis the Lanczos coefficients $\{b_{n}\}$ are replaced by $\sqrt{b_{n}c_{n}}$ and are related to the moments $\{\mu_{n}\}$ as NSSrivatsa:2023pby :
-$b_{n}c_{n}$ $=\frac{\tau_{n+1}(t)\tau_{n-1}(t)}{\tau_{n}^{2}(t)},\quad n\geq 0.$ (12)
+For general cases, the transformation symmetrizes the off-diagonals, allowing for a form of a generalized moments method for Krylov complexity, which accounts for dissipative effects, where the imaginary parts of $a_{n}$ typically suppress complexity growth.
+As mentioned above, in this transformed basis, the Lanczos coefficients $\{b_{n}\}$ are replaced by $\tilde{b}_{n}=\sqrt{b_{n}c_{n}}$ and are related to the moments $\{\mu_{n}\}$ as NSSrivatsa:2023pby:
+$\tilde{b}^{2}_{n}=b_{n}c_{n}$ $=\frac{\tau_{n+1}(t)\tau_{n-1}(t)}{\tau_{n}^{2}(t)},\quad n\geq 0\,.$ (20)
 In the following, we first consider a damped harmonic oscillator, followed by the well-known Caldeira–Leggett model as our representative open quantum systems.
-We then apply the moments method to compute and study the evolution of Krylov complexity.
+For both models, the Krylov basis is not large, and the $b_{n}$ Lanczos coefficients truncate fairly quickly, as is the case for the Harmonic Oscillator (shown in Appendix (A)).
+The Hilbert space of the system is infinite-dimensional, but because of the choice of the initial operator and associated algebra, the Lanczos $b_{n}$ coefficients terminate rapidly, as we will show later in Fig.
+(5).
+We then apply the moments method as described above to compute and study the evolution of Krylov complexity.
 3 Damped Harmonic Oscillator
@@
 The adjoint master equation, describing operator evolution, is:
-$\frac{dA}{dt}=i[H,A]+\sum_{i}\left(L_{i}^{\dagger}AL_{i}-\frac{1}{2}L_{i}^{\dagger}L_{i}A-\frac{1}{2}AL_{i}^{\dagger}L_{i}\right)\,.$ (15)
+$\frac{dO(t)}{dt}=i[H,O(t)]+\sum_{i}\left(L_{i}^{\dagger}O(t)L_{i}-\frac{1}{2}L_{i}^{\dagger}L_{i}O(t)-\frac{1}{2}O(t)L_{i}^{\dagger}L_{i}\right)\,.$ (23)
 The first term generates unitary evolution, while the remaining terms account for transitions due to interactions with the bath.
-The operators $L_{i}$ are jump operators, with $L_{i}^{\dagger}AL_{i}$ driving quantum transitions and the terms $\frac{1}{2}L_{i}^{\dagger}L_{i}A$ and $\frac{1}{2}AL_{i}^{\dagger}L_{i}$ correcting for cases where no jump occurs Preskill1998 .
+The operators $L_{i}$ are jump operators, with $L_{i}^{\dagger}AL_{i}$ driving quantum transitions and the terms $\frac{1}{2}L_{i}^{\dagger}L_{i}O(t)$ and $\frac{1}{2}O(t)L_{i}^{\dagger}L_{i}$ correcting for cases where no jump occurs Preskill1998.
 Using the framework mentioned in Sec.
@@
 Unlike the harmonic oscillator, which is a closed system, the damped oscillator loses information to the environment without any backflow.
+This is captured in the presence of the complex $a_{n}$ Lanczos coefficients, as for the harmonic oscillator and the damped oscillator there is only one non-zero $b_{n}$ coefficient.
 For the harmonic oscillator, the complexity displays periodic oscillations between zero and one, reflecting the evolution of the position operator into the momentum operator and vice versa, as dictated by the canonical commutation relations.
@@
 This interpretation is further supported by the probability amplitude plots as shown in Fig.
-(2): in the underdamped case, the slower decay indicates a less rapid loss of information to the bath.
-4 Caldeira-Leggett Model
+(3).
+For the underdamped case, the slower decay indicates a less rapid loss of information to the bath.
+4 Caldeira-Leggett master equation
 The Caldeira-Leggett model is a system-bath model designed to study quantum Brownian motion CALDEIRA1983374; Caldeira:1982iu; PhysRevLett.46.211.
@@
 $H=\frac{p^{2}}{2m}+\frac{m\omega^{2}x^{2}}{2}+\sum_{n}\left(\frac{p_{n}^{2}}{2m_{n}}+\frac{m_{n}\omega_{n}^{2}x_{n}^{2}}{2}\right)+x\sum_{n}\kappa_{n}x_{n}\,.$ (32)
-In the Schrödinger picture, the dynamics of the reduced system are described by the Caldeira-Leggett master equation BRE02 :
-$\frac{d}{dt}\rho_{s}(t)=-\frac{i}{\hbar}[H_{s},\rho_{s}(t)]-\frac{i\gamma}{\hbar}[x,\{p,\rho_{s}(t)\}]-\frac{2m\gamma k_{B}T}{\hbar^{2}}[x,[x,\rho_{s}(t)]],$ (25)
-where $H_{s}=\frac{p^{2}}{2m}+\frac{m\omega^{2}x^{2}}{2}$.
-The first term describes coherent dynamics, the second term, proportional to the relaxation rate $\gamma$, drives dissipation by damping momentum expectation values, and the third term, proportional to temperature $T$, induces decoherence through thermal fluctuations Schlosshauer_2019 ; BRE02 .
+In the Schrödinger picture, the dynamics of the reduced system are described by the Caldeira-Leggett master equation:
+$\frac{d}{dt}\rho_{s}(t)=-\frac{i}{\hbar}[H_{s},\rho_{s}(t)]-\frac{i\gamma}{\hbar}[x,\{p,\rho_{s}(t)\}]-\frac{2m\gamma k_{B}T}{\hbar^{2}}[x,[x,\rho_{s}(t)]]$ (33)
+where once again $H_{s}$ is the standard Hamiltonian of the quantum harmonic oscillator.
+As seen from the derivation the terms in the master equation come directly from the dissipative and noise kernels respectively, as such the first term describes coherent dynamics, the second term, proportional to the relaxation rate $\gamma$, drives dissipation by damping momentum expectation values, and the third term, proportional to temperature $T$, induces decoherence through thermal fluctuations Schlosshauer_2019; BRE02.
 The Caldeira-Leggett master equation is not directly in Lindblad form, as its dissipator is not positive definite.
-In the high-temperature limit, this can be remedied by adding a minimally invasive term $\left(-\frac{\gamma}{8mk_{B}T}[p,[p,\rho_{s}(t)]]\right)$ Ramazanoglu_2009 ; BRE02 ; Schlosshauer_2019 .
+In the high-temperature limit, this can be remedied by adding a minimally invasive term, as was seen in the discussion in the derivation of the C-L master equation Ramazanoglu_2009; BRE02; Schlosshauer_2019.
 The Lindbladian form facilitates analysis of operator dynamics in the adjoint form:
-$\frac{d}{dt}A(t)=i\left[H_{s}+\frac{\gamma}{2}(xp+px),A\right]+\gamma\left(L^{\dagger}AL-\frac{1}{2}L^{\dagger}LA-\frac{1}{2}AL^{\dagger}L\right),$ (26)
+$\frac{d}{dt}O(t)=i\left[H_{s}+\frac{\gamma}{2}(xp+px),O(t)\right]+\gamma\left(L^{\dagger}O(t)L-\frac{1}{2}L^{\dagger}LO(t)-\frac{1}{2}O(t)L^{\dagger}L\right),$ (34)
 with the Lindblad jump operator defined as
 $L=Ax+iBp$ (35)
-where $A=\sqrt{\frac{4mk_{B}T}{\hbar^{2}}}$ and $B=\sqrt{\frac{1}{4mk_{B}T}}$. The approximate time evolution for an operator $O(t)$ is:
+where the coefficients $A$ and $B$ are the following:
+$A=\sqrt{\frac{4mk_{B}T}{\hbar^{2}}}\,,\quad B=\sqrt{\frac{1}{4mk_{B}T}}\,.$ (36)
+In a similar manner to how we computed the operator evolution in the case of the damped oscillator.
+We begin with an initial operator $O(0)=p^{2}$, then the commutator structure of the RHS of the master equation (34) yields the following operator expression  [^1] The proportionality sign is there to indicate that the RHS of this expression does not have any time-dependent contribution.:
 $\frac{d}{dt}O(t)$ $\propto-\omega^{2}(xp+px)-2\gamma p^{2}+\gamma(A^{2}-2ABp^{2})\,$, (37)
-which in terms of creation and annihilation basis can be written as
-$\frac{d}{dt}O$ $=-\omega^{2}(xp+px)+(-2\gamma-2\gamma AB)p^{2}+\gamma A^{2}(aa^{\dagger}-a^{\dagger}a).$ (29)
-Here the proportionality factor is taken to be unity.
-Factorizing the time dependence of the creation and annihilation operators as $a(t)=f^{*}(t)a$ and $a^{\dagger}(t)=f(t)a^{\dagger}$, and solving the differential equation we obtain the following time evolved operator
+Gathering terms and writing the identity term in terms of creation and annihilation operators, the expression can be written as
+$\frac{d}{dt}O(t)$ $\propto-\omega^{2}(xp+px)+(-2\gamma-2\gamma AB)p^{2}+\gamma A^{2}(aa^{\dagger}-a^{\dagger}a)\,.$ (38)
+Details of the derivation are given in Appendix (B). Solving the differential equation and gathering the terms together, we obtain the following time-evolved operator
 $O(t)$ $=-\frac{\omega}{2}\Big(e^{(2i\omega-2\gamma-2\gamma AB)t}aa-e^{(\gamma A^{2}-2\gamma-2\gamma AB)t}aa^{\dagger}$
@@
 where the coefficients $C_{1},C_{2}$ and $C_{3}$ are given as,
-$C_{1}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}e^{-\beta(n+1/2)\omega}=\frac{(-1+e^{\beta\omega})^{2}}{3(e^{\beta\omega}+1)^{2}}\,$, (33)
+$C_{1}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}e^{-\beta(n+1/2)\omega}=\frac{(-1+e^{\beta\omega})^{2}}{3(e^{\beta\omega}+1)^{2}}\,$,
 $C_{2}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}ne^{-\beta(n+1/2)\omega}=\frac{-1+e^{\beta\omega}}{3(e^{\beta\omega}+1)^{2}}\,$, (42)
-$C_{3}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}n^{2}e^{-\beta(n+1/2)\omega}=\frac{1}{3+3e^{\beta\omega}}\,.$ (35)
+$C_{3}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}n^{2}e^{-\beta(n+1/2)\omega}=\frac{1}{3+3e^{\beta\omega}}\,.$
 These coefficients and the two-point function are evaluated at various temperatures, consistent with the high-temperature limit of the Caldeira–Leggett master equation.
@@
 $H_{I}(t)=e^{iH_{0}t}H_{I}e^{-iH_{0}t}=\sum_{n}|n\rangle\langle n|\otimes B_{n}(t)\,.$ (46)
-As a result of the commutation relation the basis states $|n\rangle$ are not effected by the coupled dynamics. This gives us the following system bath state:
+As a result of the commutation relation, the basis states $|n\rangle$ are not affected by the coupled dynamics. This gives us the following system bath state:
 $|\Psi(t)\rangle=\sum_{n}c_{n}|n\rangle\otimes|\phi_{n}(t)\rangle\,.$ (47)
@@
 Due to the interaction, the reservoir carries information on the system state.
-These coherences/correlations are still present in the system’s reduced density matrix:
+These coherence/correlations are still present in the system’s reduced density matrix:
 $\rho_{s}(t)=\mathrm{Tr}_{B}{|\Psi(t)\rangle\langle\Psi(t)|}=\sum_{n}\sum_{m}c_{n}c^{*}_{m}|n\rangle\langle m|\langle\phi_{m}(t)|\phi_{n}(t)\rangle\,.$ (48)
@@
 […]
-The quantity $\Gamma_{nm}(t)$ characterizes the behaviour of the off diagonal terms in the reduced density matrix and is known as the decoherence function.
+The quantity $\Gamma_{nm}(t)$ characterizes the behaviour of the off-diagonal terms in the reduced density matrix and is known as the decoherence function.
 For the Caldeira-Leggett model, the decoherence function is BRE02:
-$\Gamma(t)$ $=-\frac{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}{8\sigma_{0}^{2}}\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}},$ (43)
-$\tilde{\Gamma}(t)$ $=\frac{8\sigma_{0}^{2}\Gamma(t)}{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}=\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}},$ (44)
+$\Gamma(t)$ $=-\frac{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}{8\sigma_{0}^{2}}\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}}\,$,
+$\tilde{\Gamma}(t)$ $=\frac{8\sigma_{0}^{2}\Gamma(t)}{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}=\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}}$ (50)
 where $I$, $G_{1}$, and $G_{2}$ are the inhomogeneous and homogeneous solutions of the operator $x(t)$ evolution, and $\Delta x$ and $\Delta p$ are the initial separations of superposed Gaussian wave packets in position and momentum space, respectively.
@@
 (4), it is clear that the decoherence rate saturates, as described in Schlosshauer_2019.
-This occurs when $x-x^{\prime}$ reaches the maximum coherence length of the oscillator environment.
-This length is the characteristic scale over which the phase coherences of the environment, and hence the bath correlations, remain significant.
+This occurs when the separation between two system positions $x$ and $x^{\prime}$ reaches the maximum coherence length of the environment [^2] When the separation between $x$ and $x^{\prime}$ exceeds the coherence length, as the environment is performing a type of “measurement” on the system, it cannot distinguish any further between the points causing decoherence to saturate.
 The saturation of the decoherence function implies that the off-diagonal elements of the density matrix vanish.
@@
 We therefore suspect that Krylov complexity may be insensitive to decoherence, a possibility we will explore in the next subsection.
-4.2 Krylov Complexity
+4.2 Lanczos Coefficients and Krylov Complexity
+Now we will turn our attention to the Krylov complexity.
+Before presenting the results for it, we will first examine the nature of the Lanczos coefficients, which subsequently enter into the computation of the complexity.
+Note that the dimension of the Krylov space largely depends on the underlying algebra and the choice of the initial operator.
+For the cases considered in this paper, we start with a simple operator to probe the effects of dissipation and decoherence in these models.
+Below is a plot of the Lanczos coefficients $\tilde{b}_{n}$ vs n.n\,.
+Figure 5: $\tilde{b}_{n}$ Lanczos coefficients for the C-L model, with $\omega=1$ and $\gamma=1/10$.
+Similar to the results for the closed system (the quantum harmonic oscillator), and as expected, the Lanczos coefficients terminate quickly as a result of the algebra and initial operator choice.
+While the $a_{n}$ coefficients are complex and account for the dissipative features in the Krylov complexity, the $\tilde{b}_{n}^{\prime}s$ remain real.
+It is clear that, simply by examining the behavior of the Lanczos coefficients, we cannot obtain any significant information about the dynamics, especially regarding the role of dissipation.
+Hence we turn our attention to the Krylov complexity.
 As outlined in Sec.
@@
 4.3 Decomposing the Master Equation into Individual Terms
-Analzing the master equation it is clear that under particular operator choices and assumptions we can isolate the contributions from individual terms in the master equation and study how these terms contribute to Krylov complexity.
-Therefore in the following sections we will consider particular choices that allows us to investigate the individual contributions of decoherence and dissipation.
+Analyzing the master equation, it is clear that under particular operator choices and assumptions, we can isolate the contributions from individual terms in the master equation and study how these terms contribute to Krylov complexity.
+Therefore, in the following sections, we will consider particular choices that allow us to investigate the individual contributions of decoherence and dissipation.
 Decoupling Dissipation from the Master Equation
@@
 This yields a master equation that accounts only for the coherent system dynamics and the environment-induced decoherence, as discussed in Ferialdi_2017; BRE02; Schlosshauer_2019.
-Figure 6: Left: Time evolution of Krylov complexity (K-complexity) for $\omega=1$ and $\gamma=1/100$.
+Using the same operator choice as in Sec.
+(4), we compute the two-point correlation function:
+$\langle O(t)O\rangle_{\beta}$ $=e^{-2i\omega t}(C_{3}+3C_{2}+2C_{1})+e^{(-8m\gamma k_{B}T/\hbar^{2})t}(2C_{3}+3C_{2}+C_{1})$
+$\quad+e^{(8m\gamma k_{B}T/\hbar^{2})t}(2C_{3}+C_{2})+e^{2i\omega t}(C_{3}-C_{2})$ (52)
+Using the correlation function, we compute the K-complexity, which only consists of the coherent system dynamics and thermal decoherence term, as is displayed in Fig.(8) below:
+Figure 8: Left: Time evolution of Krylov complexity (K-complexity) for $\omega=1$ and $\gamma=1/10$.
 Right: Time evolution of Krylov complexity for the same parameters, with the decoherence function $\tilde{\Gamma}(t)$ overlaid in orange.
@@
 For $\rho_{s}(t)$ in the form given by (48), it is evident that at least one of the diagonal elements of the reduced density matrix is time dependent, while at least one of the off-diagonal elements contains a term without time dependence.
-Consequently, there is no clear separation of the correlations carried by the off-diagonal terms, and the Krylov complexity cannot be reliably used to indicate the decay of coherences.
+Consequently, there is no clear separation of the correlations carried by the off-diagonal terms, and the Krylov complexity cannot be reliably used to indicate the decay of coherence.
 Finally, the inclusion of the dissipative term determines how rapidly the complexity decays or grows.
@@
 In this work, we investigate Krylov complexity in open quantum systems, focusing on the damped harmonic oscillator and the Caldeira–Leggett model.
+We emphasize that the choice of relatively simple models is deliberate, providing the analytical tractability (especially for evaluating correlation functions) essential to dissect the roles of dissipation and decoherence via Krylov techniques.
+Although Krylov complexity and associated Lanczos coefficients have traditionally been leveraged to explore quantum chaos, this study re-purposes them to pinpoint these effects in open quantum systems.
+As far as we are aware, ours is an early effort to deploy Krylov methods for this purpose, shedding new light on the intricate interplay between dissipation and decoherence.
 Our results show that Krylov complexity effectively captures dissipative and decoherence effects, revealing distinct behaviors in these systems.
@@
 $[H,O]$ $=\omega\alpha(a^{{\dagger}}-a)\,$, (56)
-$[H,[H,O]]$ $=\omega^{2}\alpha(a^{{\dagger}}+a)$ (50)
-$,,$ (51)
+$[H,[H,O]]$ $=\omega^{2}\alpha(a^{{\dagger}}+a)\,$, (57)
 $[H,[H,[H,O]]]$ $=\omega^{3}\alpha(a^{{\dagger}}-a)$ (58)
@@
 Figure 10: Time evolution of Krylov complexity for Harmonic oscillator with $\omega=1$ and $m=1$.
+Appendix B Correlation functions and Time evolved operators
+Here we show the general procedure followed to compute the time-evolved operators for the Caldeira-Leggett master equation in Sec.
+(4).
+We only show the procedure followed in the first case, since for the rest of the operators, only a slight change is needed.
+We begin with the full form of the Lindblad master equation as given in (34), with the initial operator being $O(0)=p^{2}$, this is because it is the simplest form of operator that allows us to obtain contributions from both the dissipative and decoherence terms.
+The Lindblad jump operator $L$ is given in (35).
+The general procedure we will follow here is similar to Sec.
+(3).
+The evolution of such operators when written in terms of creation and annihilation operators can, in general, be written in terms of some time-dependent functions $f_{i}(t)\in\mathbb{C}$.
+As such, initially, we will leave the function, as well as the time dependence, implicit, focusing only on the operator algebra, and reintroduce the time dependence later.
+We begin by considering the individual contributions from each term of (34).
+From the first term, for the coherent dynamics, we have:
+$i[H_{s},p^{2}]$ $=i\left[\left(\frac{p^{2}}{2}+\frac{\omega^{2}x^{2}}{2}\right),p^{2}\right]\,$,
+$=i\omega^{2}\left[\frac{x^{2}}{2},p^{2}\right]\,$,
+$=i\omega^{2}\left[\frac{x^{2}}{2},p\right]p+i\omega^{2}p\left[\frac{x^{2}}{2},p\right]\,$,
+$=i\omega^{2}(ixp+ipx)\,$,
+$=-\omega^{2}xp-\omega^{2}px\,.$ (64)
+The second term of (34), which behaves like a Hamiltonian contribution and comes from the dissipator, yields the following:
+$i\frac{\gamma}{2}[xp+px,p^{2}]$ $=i\frac{\gamma}{2}[xp,p^{2}]+i\frac{\gamma}{2}[px,p^{2}]\,$,
+$=i\frac{\gamma}{2}[2ip^{2}]+i\frac{\gamma}{2}[2ip^{2}]\,$,
+$=-2\gamma p^{2}\,.$ (65)
+Now we tackle the contributions which come from the dissipator in the Lindblad master equation, the third term of the master equation (34) gives us:
+$L^{\dagger}p^{2}L$ $=(Ax-iBp)(p^{2})(Ax+iBp)\,$,
+$=A^{2}xp^{2}x+iABxp^{3}+B^{2}p^{4}-iABp^{3}x\,.$ (66)
+From the fourth term (of (34)), we get:
+$-\frac{1}{2}L^{\dagger}Lp^{2}$ $=(Ax-iBp)(Ax+iBp)(p^{2})\,$,
+$=-\frac{1}{2}(A^{2}x^{2}p^{2}+iABxp^{3}-iABpxp^{2}+B^{2}p^{4})$ (67)
+and finally, from the fifth term of (34) we get:
+$-\frac{1}{2}p^{2}L^{\dagger}L$ $=(p^{2})(Ax-iBp)(Ax+iBp)\,$,
+$=-\frac{1}{2}(A^{2}p^{2}x^{2}+iABp^{2}xp-iABp^{3}x+B^{2}p^{4})\,.$ (68)
+Combining all of them, we get,
+$S$ $=\gamma(A^{2}xp^{2}x-\frac{1}{2}A^{2}x^{2}p^{2}-\frac{1}{2}A^{2}p^{2}x^{2})+\gamma(\frac{i}{2}ABxp^{3}-\frac{i}{2}ABp^{3}x)+\gamma(\frac{i}{2}ABpxp^{2}-\frac{i}{2}ABp^{2}xp)\,$,
+$=\gamma(A^{2}xp^{2}x-\frac{1}{2}A^{2}x^{2}p^{2}-\frac{1}{2}A^{2}p^{2}x^{2})+\gamma\frac{iAB}{2}[x,p^{3}]+\gamma\frac{iAB}{2}[pxp,p]\,$,
+$=\gamma(A^{2}xp^{2}x-\frac{1}{2}A^{2}x^{2}p^{2}+[x^{2},p^{2}]-\frac{1}{2}A^{2}x^{2}p^{2})+\gamma\frac{iAB}{2}[x,p^{3}]+\gamma\frac{iAB}{2}[pxp,p]\,$,
+$=\gamma(A^{2}xp^{2}x-A^{2}x^{2}p^{2}+[x^{2},p^{2}])+\gamma\frac{iAB}{2}[x,p^{3}]+\gamma\frac{iAB}{2}[pxp,p]\,$,
+$=\gamma(A^{2}x[p^{2},x]+[x^{2},p^{2}])+\gamma\frac{iAB}{2}[x,p^{3}]+\gamma\frac{iAB}{2}[pxp,p]\,$,
+$=\gamma(A^{2}-2ABp^{2})\,$, (69)
+This leaves us with the following:
+$\frac{d}{dt}O$ $\propto-\omega^{2}xp-\omega^{2}px-2\gamma p^{2}+\gamma(A^{2}-2ABp^{2})\,$,
+$\propto-\omega^{2}xp-\omega^{2}px+(-2\gamma-\gamma 2AB)p^{2}+\gamma A^{2}(aa^{\dagger}-a^{\dagger}a)\,.$ (70)
+This is the equation (38) that we have used in the main text.
+Now to solve for the $O(t),$ we first recall that $a(t)=f^{*}(t)a$ and similarly $a^{{\dagger}}(t)=f(t)a^{{\dagger}}$ as discussed in Sec.
+(3).
+We then rewrite every term of (70) in terms of $a(t)$ and $a^{{\dagger}}(t)$ in the following way,
+$(-\omega^{2})xp$ $=\underbrace{(-\frac{i}{2}\omega^{2})}_{M}(a^{\dagger}(t)a(t)-a^{\dagger}(t)a^{\dagger}(t)+a(t)a(t)-a(t)a(t)^{\dagger})\,$,
+$(-\omega^{2})p\,x$ $=\underbrace{(-\frac{i}{2}\omega^{2})}_{N}(a(t)a^{\dagger}(t)-a^{\dagger}(t)a^{\dagger}(t)+a(t)a(t)-a^{\dagger}(t)a(t))\,$,
+$(-2\gamma-2\gamma AB)p^{2}$ $=\underbrace{\frac{\omega}{2}(2\gamma+2\gamma AB)}_{O}(a(t)a(t)-a(t)a^{\dagger}(t)-a^{\dagger}(t)a(t)+a^{\dagger}(t)a^{\dagger}(t))\,.$ (71)
+We now have the following form of the differential equation.
+$\frac{d}{dt}O$ $=-(-M+N+O)f_{1}(t)a^{\dagger}a+(-M-N-O)f_{2}(t)a^{\dagger}a^{\dagger}$
+$+(M+N+O)f_{3}(t)aa-(M-N+O)f_{4}(t)aa^{\dagger}\,.$ (72)
+The functions $f_{1},f_{2},f_{3},f_{4}$ can be read off from (28). Finally, we get an explicit form of the time-evolved operator as:
+$O(t)$ $=-\frac{\omega}{2}(\exp[(2i\omega-2\gamma-2\gamma AB)t]aa$
+$-\exp[(\gamma A^{2}-2\gamma-2\gamma AB)t]aa^{\dagger}$
+$-\exp[(-\gamma A^{2}-2\gamma-2\gamma AB)t]a^{\dagger}a$
+$+\exp[(-2i\omega-2\gamma-2\gamma AB)t]a^{\dagger}a^{\dagger})\,.$ (73)
+Equipped with this, we can therefore compute the correlator:
+$\langle O(t)O\rangle_{\beta}$ $=\frac{\mathrm{Tr}\{\rho_{\beta}O(t)O\}}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\,$,
+$=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}e^{-\beta(n+1/2)\omega}\langle n|\frac{\omega^{2}}{4}(e^{(2i\omega-2\gamma-2\gamma AB)t}aa-e^{(\gamma A^{2}-2\gamma-2\gamma AB)t}aa^{\dagger}$
+$-e^{(-\gamma A^{2}-2\gamma-2\gamma AB)t}a^{\dagger}a+e^{(-2i\omega-2\gamma-2\gamma AB)t}a^{\dagger}a^{\dagger})(a^{{\dagger}}a^{{\dagger}}-a^{{\dagger}}a-aa^{{\dagger}}+aa)|n\rangle\,$,
+$=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\frac{\omega^{2}}{4}\sum_{n}e^{-\beta(n+1/2)\omega}(e^{(2i\omega-2\gamma-2\gamma AB)t}(n^{2}+3n+2)$
+$+e^{(\gamma A^{2}-2\gamma-2\gamma AB)t}(2n^{2}+3n+1)+e^{(-\gamma A^{2}-2\gamma-2\gamma AB)t}(2n^{2}+n)$
+$+e^{(-2i\omega-2\gamma-2\gamma AB)t}(n^{2}-n))\,.$ (74)
+Here, the choice of normalization is the common choice for normalization to ensure that at C(0)=1.C(0)=1\,. Now we use the following to perform the sum over $n$,
+$\begin{split}&\sum_{n}\langle n|a^{{\dagger}}a^{{\dagger}}aa|n\rangle=n^{2}-n\,,\quad\sum_{n}\langle n|aaa^{{\dagger}}a^{{\dagger}}|n\rangle=n^{2}+3n+2\,,\\ &\sum_{n}\langle n|aa^{{\dagger}}aa^{{\dagger}}|n\rangle=n^{2}+2n+1\,,\quad\sum_{n}\langle n|aa^{{\dagger}}a^{{\dagger}}a|n\rangle=n^{2}+n\,,\\ &\sum_{n}\langle n|a^{{\dagger}}aa^{{\dagger}}a|n\rangle=n^{2}\,,\quad\sum_{n}\langle n|a^{{\dagger}}aaa^{{\dagger}}|n\rangle=n^{2}+n\,.\end{split}$
+Finally, one can use the expressions for $C_{1},C_{2},C_{3}$ as mentioned in (42) of Sec .
+4 to bring it into a more compact form.
+The procedure outlined above was used in computing the operators $O(t)$ and correlators for the entirety of Sec.
+(4).